### v1

# Additive Sink Subtraction

Abstract
Subtraction games are a classical topic in Combinatorial Game Theory. A result of Golomb (1966) shows that every subtraction game with a finite move set has an eventually periodic nim-sequence, but the known proof yields only an exponential upper bound on the period length. Flammenkamp (1996) conjectures a striking classification for three-move subtraction games: non-additive rulesets exhibit linear period lengths of the form “the sum of two moves”, where the choice of which two moves displays fractal-like behavior, while additive sets $S=\{a,b,a+b\}$ have purely periodic outcomes with linear or quadratic period lengths. Despite early attention in Winning Ways (1982), the general additive case remains open. We introduce and analyze a dual winning convention, which we call sink subtraction. Unlike the standard wall convention, where moves to negative positions are forbidden, the sink convention declares a player the winner upon moving to a non-positive position. We show that additive sink subtraction admits a complete solution: the nim-sequence is purely periodic with an explicit linear or quadratic period formula, and we conjecture a duality between additive sink subtraction and classical wall subtraction. Keywords: Additive Subtraction Game, Nimber, Periodicity, Sink Convention.
1 Introduction
We study classical combinatorial subtraction games [1], with a twist on the terminal condition. Let $S\subset\mathbb{N}=\{1,2,\ldots\}$. For any $x\in\mathbb{N}$ and all $s\in S$, $x-s\in\mathbb{Z}$ is a move option. The sink is an infinite sequence of terminal P-positions for all non-positive integers. No move is possible from a position in the sink. We study two-player impartial games in the normal-play convention, so “a player who cannot move loses”, and this is the same as “a player who plays into the sink wins”. We denote this ruleset family by sink subtraction. Observe that, if $S=\mathbb{N}$, we play the game of nim and sink subtraction coincides with wall subtraction (our terminology), where all negative heap sizes are forbidden. In Figure 1, we illustrate with the ruleset $S=\{1,5\}$ how sink vs. wall play can differ.
Figure 1: sink and wall subtraction, respectively. Suppose the ruleset is $S=\{1,5\}$. To the Left, the boy can choose to move into the sink or he can play to the tip of the jetty, and let daddy win. To the Right, the boy has only one move, because the larger move is prohibited by the wall. Thus the wall-game is determined already by parity.
Remark 1.
In the classical normal-play setting, we may imagine that a concrete wall is placed at position 1-1, preventing players from entering any negative number. Curiously, if we force negative positions to be terminal N-positions, information is lacking about how to play a disjunctive sum of games; we need to assign non-zero nimbers to every nonpositive position. Hence, in this sense, the standard wall convention actually represents infinitely many possible termination rules, while the sink convention yields a unique interpretation (since all P-positions have nim-value zero).
Indeed, if $S$ is finite, the two ruleset families sink and wall usually satisfy distinct nimber patterns. A game’s nimber (a.k.a. nim-value) is defined via the standard minimum exclusive rule; for any finite $X\subset\mathbb{N}_{0}$, let $\mathrm{mex}X=\min\{\mathbb{N}_{0}\setminus X\}$, where $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. Then the nim value algorithm of $x$ is $v(x)=\mathrm{mex}\{v(x^{\prime})\}$, where $x^{\prime}$ ranges over the options of $x$. Note that this gives value zero for all positions in the sink.
For example, if $S=\{2,5\}$, then starting from $x=0$, ‘wall’ has a purely periodic nim-sequence of period length 7,
${\bf 0,0,1,1,0,2,1},0,0,1,1,0,2,1\ldots,$
while ‘sink’, starting from $x=1$, begins
$1,1,2,{\bf 2,1,0,0,1,1,0},2,1,0,0,1,1,\ldots$
with a pre-period of size three but the same period length 7. The periodic patterns are the same but occur in a different order: 0011021 and 2100110 respectively. To the best of our knowledge, sink subtraction has not yet been studied in the literature. It is the purpose of this note to enter this unpaved territory. We use the convention of $x=0$ belonging to the sink, so we start nim-sequences by the (non-zero) value of position one.
Of course, the folklore periodicity theorem holds for rulesets where the terminal condition is a sink instead of a ‘wall’, with exactly the same proof.
Theorem 2 (Folklore [3]).
Any instance of sink subtraction has an ultimately periodic nim-sequence.
Proof.
Let $S$ be the subtraction set of sink subtraction. The window of options for any position in a subtraction game is determined by $\max S$. The window symbols are nimbers, computed recursively. By the mex-rule, the number of nimbers is bounded by $\rho=|S|$.
Thus, the number of possible combinations of symbols in a window of size $\max S$ is $\rho^{\max S}$. This number is finite as $\max S$ is finite.
The window of options of a position $x$ is $\{x-\max S,\ldots,x-1\}$, while for the position $x+1$, the window of options is $\{x+1-\max S,\ldots,x\}$, and so on, thus making the number of windows of options infinite.
The number of combinations of symbols is finite, while the number of windows of options is infinite; this implies that some window symbol combination will repeat. The first occurrence of such a repetition will define the period length. It is possible that, for an initial sequence, window symbol combinations do not repeat; if so, this forms the pre-period. ∎
Note that Theorem 2 establishes eventual periodicity for any terminating condition; in general, one can assign arbitrary patterns of forced nimbers for negative integers, and the folklore theorem still holds. There is a recent paper that discusses such settings [miklós2023superpolynomial], and they establish super-polynomial period length for three-move games with specific such relaxed winning conditions.
The purpose of this note is to establish the period lengths of the ruleset family additive sink subtraction $S$ with subtraction set
$S(m,\delta)=\{m,m+\delta,2m+\delta\},$ (1)
where $m,\delta\in\mathbb{N}$. There is a conjecture from Flammenkamp’s PhD thesis [2] concerning the periodicity of Additive Wall Subtraction, and this is to some degree paraphrasing the treatment in Winning Ways [1], where the linear-period-length part of the conjecture has been proved. As far as the authors know, the non-linear (quadratic) part has not yet been established. [^1] Non-additive three-move games are conjectured to have period lengths of the form “the sum of two moves”, where “which two moves?” appear to satisfy some fractal patterns [2]. Let us state our main result.
Theorem 3 (Main Theorem).
Let $m,\delta\in\mathbb{N}$. Consider additive sink subtraction with subtraction set $S(m,\delta)=\{m,m+\delta,2m+\delta\}$, and let $d:=\delta\ \pmod{2m}$, with $0\leq d<2m$. Then $S$ is purely periodic, and its period length is
$p(m,\delta)=\begin{cases}3m+2\delta-d,&\mathrm{if}\ d\in[0,m];\\ m(m+2\delta+d)/\gcd(m,d),&\mathrm{otherwise}.\end{cases}$
2 A proof of the main result
The rest of the paper concerns the proof of Theorem 3. We distinguish two cases:
  1. (i)
    $0\leq d\leq m$;
  2. (ii)
    $m<d<2m$.
Note that, if $m=1$, only item (i) applies. Let us prove item (i). Write the moves as
$s_{1}:=m,\qquad s_{2}:=m+\delta,\qquad s_{3}:=2m+\delta.$
Define the candidate period as the word on the alphabet $\{0,1,2,3\}$:
$(1^{m}2^{m})^{a}\,3^{d}\,0^{m}\,(3^{m}0^{m})^{a-1},$ (2)
where $a:=\frac{\delta-d}{2m}+1$. The length of the candidate period (2) is
$2ma+d+m+2m(a-1)$ $=2m\!\left(\frac{\delta-d}{2m}+1\right)+d+m+2m\!\left(\frac{\delta-d}{2m}\right)$
$=\delta-d+2m+d+m+\delta-d$
$=3m+2\delta-d,$
as claimed.
Let us verify that the nim-sequence starting at $x=1$ follows the pattern of the candidate period (2).
Case 1: the initial word $(1^{m}2^{m})^{a}$. For $1\leq x\leq 2ma=\delta-d+2m\leq s_{3}$, we have $x-s_{3}\leq 0$, and hence, by the definition of the sink, $v(x-s_{3})=0$. Since $s_{1}=m$, the mex rule forces alternating blocks of 11 and $2$, each of length $m$. The move $s_{2}=m+\delta$ does not interfere: the largest relevant options lie in
[ $2ma-m-s_{2},\;2ma-s_{2}\,]=[-d,\;m-d]\subseteq(-\infty,m],$
whose nim-values are at most $1$. Thus $v(x)=2$ at the appropriate positions, completing the first part.
Case 2: the $3^{d}$ factor. For $2ma<x\leq 2ma+d$, we have
$v(x-s_{1})=2,\quad v(x-s_{2})=1,\quad v(x-s_{3})=0,$
since $x-s_{3}\leq 0$. Hence $v(x)=3$ for exactly $d$ consecutive positions.
Case 3: the $0^{m}$ factor. For $2ma+d<x\leq 2ma+d+m=3m+\delta$, all options satisfy $x-s_{3}>0$, so no option has value 0. Thus the mex rule produces $m$ consecutive zeros.
Case 4: the trailing $(3^{m}0^{m})^{a-1}$. For the remaining 2m(a1)2m(a-1) positions, the same mechanism repeats. The moves $s_{3}$ and $s_{2}$ together generate options of values 11 and $2$, while $s_{1}=m$ toggles between producing and avoiding the value 0. Since $s_{3}-s_{2}=m$, it suffices to note that for all such $x$, both $x-s_{3}$ and $x-s_{2}$ lie in $[1,2ma]$, where the values are already known. This forces alternating blocks of $3^{m}$ and $0^{m}$.
Finally, observe that the next $2m+\delta$ positions reproduce the initial segment $(1^{m}2^{m})^{a}3^{d}$. Hence the nim-sequence is purely periodic with period length $p(m,\delta)=3m+2\delta-d$.
Consider item (ii). For convenience we restrict $m<\delta<2m$, and let $\delta=m+k$. (The general case turns out to be an immediate consequence.) As an example, let us begin by describing the periodic word for any rulesets of the form “the lower case”, i.e. whenever $k=1$. The word is described by consecutive “$B$-blocks” of the form
$1^{m-i}0^{i}2^{m-i}1^{1+i}3^{m-1-i}2^{1+i}0^{m-i}3^{1+i},$
for $0\leq i\leq m-1$, and at last, we concatenate the factor $0^{m}$. By increasing the ‘block speed’, the factor $1\leq k<m$, another type of blocks appears, the “$C$-blocks”, and the extreme “upper case” will favor only the $C$-blocks. The challenge is to understand the intermediate cases, with mixing between the block types; in Figure 2, we illustrate all periods for the case of $m=19$, $0<k<19$.
Figure 2: Nimber periods for $m=19$ and $0<k<19$, starting from $x=1$. The colors correspond to the nimbers; purple (0), blue (1), green (2) and yellow (3). Zoom in for details.
Let $x]_{y}$ denote the smallest nonnegative representative of $x\pmod{y}$. Let $[x]$ denote the smallest nonnegative representative of $x\pmod{m}$. Let $x]^{+}$ denote the smallest positive representative of $x\pmod{m}$. For given $m,k$ and $i$, let $\alpha(i)=[ki]$, let $\beta(i)=[k(i+1)]$ and let $\gamma(i)=[k(i+1)]^{+}$. We drop $i$ in these notations, when the context allows.
For given $1\leq k<m$, $\delta=m+k$, for $0\leq i\leq m-1$, define the prefix word
$A(i)=1^{m-\alpha}0^{\alpha}2^{m-\alpha}.$ (3)
If
$\gamma>\alpha,$ (4)
define the $i$th block as
$B(i)=A(i)1^{\gamma}3^{m-\gamma}2^{\gamma}0^{m-\alpha}3^{\gamma},$ (5)
and otherwise, if (4) does not hold, define the $i$th block as
$C(i)=A(i)1^{m}2^{m}1^{\beta}0^{k-\beta}3^{m-k}2^{\beta}3^{k-\beta}0^{m}3^{\beta}.$ (6)
Observe that the $C$-block in (6) has two exponents of the form $k-\beta$. Due to the reverse of inequality (4), for a given $i$, $\alpha>\gamma\geq\beta$, then, by modular arithmetic, we must have $k>\beta$, so this exponent is always positive.
For given $m$ and $k$, define the bijection $\varphi:[0,m-1]\rightarrow B^{*}$ induced by our construction, where $B^{*}$ is the set of all $B$- and $C$-blocks. Let
$P(m,k)=\prod_{0\leq i<m}\varphi(i)\,\zeta(i),$ (7)
where $\zeta(i)=0^{m}$ if $\beta(i)=0$, and otherwise $\zeta(i)=\varnothing$. Note that $\beta=0$ whenever i=m1i=m-1, so $P$ always ends in the factor $0^{m}$. Indeed, if $m$ is prime, then $\zeta(i)=\varnothing$ if and only if i<m1i<m-1. Otherwise, we have additional inserted factors $0^{m}$, which are always preceded by a $B$-block, since $\beta=0$ implies $\gamma=m>\alpha$.
We verify that the nimbers in the additive subtraction game $S(m,2m+k,3m+k)$ are purely periodic and the (not necessarily the shortest) period is $P=P(m,k)$, if $1<k<m$. That is, if $W=(w_{i})_{i\geq 1}$ is the infinite word on the same alphabet such that, for all $i\geq 1$, $w_{i}$ is the nimber of $S$, then $W=PP\ldots$.
Observation 4.
In spite of the periodicity hinging on a quadratic expression in $m$, we are able to verify correctness in constant time. Since each block consists of a constant number of factors, 8 and 12 respectively, and the number of interactions between blocks is also a constant, the task of testing the two mex properties (anti-collision and reachability) reduces to a case by case analysis. The number of tests is still large and a computer assisted proof seems like a convenient way forward. We have instead chosen to expose the data in a number of tables, so that every entry can be verified by a human. We explain by some examples how to interpret the tables.
The relevant differences $\gamma-\alpha$ and $\beta-\alpha$ do not depend on the index $i$, and consecutive blocks trigger some convenient identities.
Lemma 5.
For given $i$,
  1. 1.
    if $\gamma>\alpha$ then $\gamma-\alpha=k>0$, and otherwise,
  2. 2.
    if $\gamma<\alpha$, then $\alpha-\beta=m-k$, $\beta<k$ and $\alpha>m-k$.
Moreover, for $i\in\{0,\ldots,m-1\},\alpha(i)=\beta(i-1)$ and for $i\in\{1,\ldots,m\},\alpha(i)=\gamma(i-1)$.
Proof.
Apply modular arithmetic. ∎
This lemma will be used often without explicit mention, as we regard these relations as basic facts that control much of the mex behavior of the rulesets.
Let us prove that the construction produces the correct period lengths. For example, if $m=5$ and $k=1$, then (by ignoring technical details) the block structure is $BBBBB0^{m}$, which satisfies the claimed period length $5(4\cdot 5+2c-2a)+5=115$. If $m=5$ and $k=4$, then the block structure is $BCCCB0^{m}$, which satisfies the claimed period length $2(4\cdot 5+8)+3(6\cdot 5+4-b+a)+5=56+99+5=160$. The cases for $m=5$ are illustrated in the upper picture of Figure 3. If $m=6$ and $k=2$, then the block structure is $BBB0^{m}BBB0^{m}$, with claimed period length $3(4\cdot 6+2c-2a)+6=90$. If $m=6$ and $k=4$, then the block structure is $BCB0^{m}BCB0^{m}$, with claimed period length $2(4\cdot 6+2c-2a)+6\cdot 6+4+b-a+6=108$. The cases for $m=6$ are illustrated in the lower picture of Figure 3.
Recall that we here assume $\delta=m+k$, $1\leq k<m$.
Figure 3: Nimber periods for $m=5,6,7$ and $0<k<m$, respectively, starting at $x=1$. The rows have different scaling as to fit the full periods in each case. In the case of $m=6$, and rows $k=2,3,4$ the smallest period is a divisor of the full row length. Hence in these cases there are more than one $\zeta$-factors. In the case of $m=7$, we have indicated with red rectangles, for $k=1$ the blocks $B(0)$ and $B(4)$; for $k=2$ the block $C(3)$; for $k=3$ the block $B(6)$; for $k=4$ the factor $\zeta(6)$; for $k=5$, the blocks $B(3)$ and $C(5)$; for $k=6$ the block $C(1)$.
Lemma 6.
For given $m,k$, we have $|P(m,k)|=m(4m+3k)/\gcd(m,k)$.
Proof.
Let $g:=\gcd(m,k)$. Write $m=gM$ and $k=gK$ with $\gcd(M,K)=1$. Recall the definitions $\alpha(i)=[ki]_{m},\beta(i)=[k(i+1)]_{m}$ and $\gamma(i)=[k(i+1)]_{m}^{+}$. Moreover, for each $i$, we build a block $B(i)$ or $C(i)$ with prefix $A(i)$, and appending an extra factor $\zeta(i)=0^{m}$ to a $B$-block if $\beta(i)=0$.
1) Periodicity in the block index $i$. For $i\geq 0$,
$\alpha(i+M)$ $=[k(i+M)]_{m}$
$=[ki+kM]_{m}$
$=[ki+(km/g)]_{m}$
$=[ki]_{m}$
$=\alpha(i),$
since $g|k$. The same calculation shows $\beta(i+M)=\beta(i)$ and hence also $\gamma(i+M)=\gamma(i)$.
Thus the whole pattern of blocks (including the presence or absence of $\zeta(i)=0^{m}$) is periodic in $i$ with period $M$. In particular, the word $P(m,k)$ is obtained by concatenating the $M$ blocks corresponding to $i=0,\dots,M-1$, repeated $g$ times. Since we are interested in the length of one period, it suffices to count the total length contributed by a single index cycle $i=0,\dots,M-1$.
2) Block lengths. From (3)–(6), the prefix $A(i)=1^{m-\alpha}0^{\alpha}2^{m-\alpha}$ has length $|A(i)|=2m-\alpha(i)$.
A direct inspection of the definitions of $B(i)$ and $C(i)$ yields:
  1. 1.
    If $\gamma(i)>\alpha(i)$ and $\beta(i)>0$ (the generic $B$-case), then $B(i)$ has length $|B(i)|=4m+2k.$
  2. 2.
    If $\gamma(i)>\alpha(i)$ and $\beta(i)=0$ (the “wrap” $B$-case with an inserted $0^{m}$), or if $\gamma(i)<\alpha(i)$ (the $C$-case), then the resulting ‘block’ (either $B(i)\zeta(i)$ or $C(i)$) has length $5m+2k.$
So every index $i$ in one cycle contributes either a short block of length $4m+2k$ or a long block of length $5m+2k$.
3) Classification via $\alpha(i)$. We claim that the type of block is determined solely by whether $\alpha(i)<m-k$ or $\alpha(i)\geq m-k$.
Indeed, if $\alpha(i)<m-k$, then $\alpha(i)+k<m$, so $k(i+1)=ki+k$ does not wrap modulo $m$, and thus
$\beta(i)=\gamma(i)=\alpha(i)+k>\alpha(i).$
Hence we are in the generic $B$-case and obtain a short block of length $4m+2k$.
Conversely, if $\alpha(i)\geq m-k$, then $\alpha(i)+k\geq m$, and so $\beta(i)$ wraps around $m$, which implies $\beta(i)=\alpha(i)+k-m<\alpha(i)$. If $\beta(i)=0$, we are in the wrapped $B$-case with an extra $\zeta(i)=0^{m}$; if $\beta(i)>0$, then $\gamma(i)=\beta(i)<\alpha(i)$ and we are in the $C$-case. In both subcases we obtain a long block of length $5m+2k$. Thus:
[…]
4) The number of short/long blocks in one cycle. Over one index cycle $i=0,\dots,M-1$, the values of $\alpha(i)$ run through the multiples of $g$ modulo $m$:
$\{\alpha(i)\mid 0\leq i\leq M-1\}=\{0,g,2g,\dots,m-g\},$
each exactly once. Write mk=g(MK)m-k=g(M-K). Then the inequality $\alpha(i)\geq m-k$ is equivalent to
$gj\geq g(M-K)\quad\Longleftrightarrow\quad j\geq M-K,$
where $\alpha(i)=gj$ with $0\leq j\leq M-1$. Hence, in one cycle:
  1. 1.
    there are exactly MKM-K indices with $\alpha(i)<m-k$ (short blocks), and
  2. 2.
    there are exactly $K$ indices with $\alpha(i)\geq m-k$ (long blocks).
5) The length of $P$. Since $P$ consists exclusively of short and long blocks, we have that
$|P(m,k)|$ $=(M-K)(4m+2k)+K(5m+2k)$
$=M(4m+2k)+Km.$
By substituting $M=m/g$ and $K=k/g$ we get
$|P(m,k)|$ $=\frac{m(4m+2k)}{g}+\frac{km}{g}$
$=\frac{m(4m+3k)}{g}$
$=\frac{m(4m+3k)}{\gcd(m,k)}.$
The blocks use two independent indices; the second one will be used throughout the tables to come.
  1. 1.
    The block index $i$ refers to the position of a block in the concatenated word $P(m,k)$. Thus $B(i)$ and $C(i)$ denote the $i$th block, of type $B$ or $C$, respectively.
  2. 2.
    Within a fixed block, we label its consecutive factors by a local index. For a $B$-block, the factors are denoted $\texttt{B1},\ldots,\texttt{B5}$, where Bj refers to the (j+3)(j+3)rd factor of that block, for $j\in\{1,\ldots,5\}$. Similarly, Aj denotes the $j$th factor of the prefix $A(i)$ associated with the same block index $i$.
Unless explicitly stated otherwise, the block index $i$ is fixed when referring to factor indices. For example, the notation B3 always refers to the sixth factor within a single $B(i)$ block, independent of the block index $i$. Whenever we refer to an object other than a block or a factor, we explicitly use the term word.
Recall that the block length of $B$ is $4m-2\alpha+2\gamma=4m+2k>3m+k$ (by $\gamma>\alpha$), while the block length of $C$ is $6m+k-\alpha+\beta=5m+2k>3m+k$ (by $\beta<\alpha$). One can check that, under block concatenation, the distance between two similar factors is always greater than $3m+k$; two A factors are similar, if they carry the same j index, and analogously for B and C.
We will treat each factor as a single entity, and use terminology such as factor-option and factor value. The advantage is that it drastically reduces the number of cases, but an apparent disadvantage is that sometimes individual positions within a factor can have options in distinct factors, for a given move. This phenomenon proves unavoidable though, and is part and parcel of the quadratic period length, as we will see. By intuition, when factor relations settle, then we approach the full period.
We use elementary algebra to justify mex-consistency. Each factor $F$, of $P$, is of the form $F=x^{y}$ with $x\in\{0,1,2,3\}$ and where $y\geq 0$ is maximum such that adjacent letters differ from $x$; the length of $F$ is $|F|=y$. The factor value of a factor $F$ is $v(F)=x$ if $F=x^{y}$. We often want to identify a factor with its corresponding set of positions. To this purpose it is convenient to temporarily identify position one with the first letter of some relevant factor, and we write F to denote the factor set, the set of such induced positions corresponding to a factor $F$. The value of a factor set is its factor value. A factor-option of $F$ is a factor $F^{\prime}$ such that there is an $x\in\texttt{F'}$ such that, for some $y\in\texttt{F}$, $x-y\in\{m,2m+k,3m+k\}$. Any such terminology is local and used as a computational device.
We display tables with factor-options and their values. Each table corresponds to a local block concatenation and verifies one of the two mex conditions. Tables 1 to 6 justify reachability, that all smaller values appear as values of options of the conjectured factor-values. Here, of course, we use standard induction principles. Tables 7 to 12 verify anti-collision, the second mex-principle, that any conjectured factor-value, cannot have the same value at the level of any factor-option.
Throughout the tables, the symbols $s_{1}$, $s_{2}$, and $s_{3}$ denote the moves $m$, $2m+k$, and $3m+k$, respectively. In all reachability tables, a dash “–” indicates that no reachability condition is required from the corresponding factor to that value. In tables for concatenations (e.g. BB), we use uppercase factor names for the current block $B(i)$ and lowercase for the preceding block B(i1)B(i-1). For example, B4 is the set of positions corresponding to the factor $0^{m-\alpha}$ in $B(i)$, while b4 is the set of positions corresponding to the factor $0^{m-\alpha^{\prime}}$ in B(i1)B(i-1); Throughout, we write $\alpha^{\prime}=\alpha(i-1)$, $\beta^{\prime}=\beta(i-1)$ and $\gamma^{\prime}=\gamma(i-1)$.
In the tables, rather than fixing a specific heap of size one, we adopt the following convention. For each row (i.e. each game value under consideration), we choose as reference point the smallest position in the leftmost factor of that value; for the sink we use position 0 as reference. Each factor is then interpreted as a set of consecutive positions, all carrying the same game value, determined relative to this reference point. In this way, the symbolic word representations, together with the specified factor lengths, are mapped unambiguously to actual game positions, allowing us to analyze move options between factors.
In the reachability tables, distances are measured from the beginning of the relevant factor to the end of the preceding block. For the factor b4, the distance to the end of the $B$-block is $d(\texttt{b4})=m-\gamma^{\prime}+\alpha^{\prime}=m-k$ (by Lemma 5), and its length is $|\texttt{b4}|=\gamma(i-1)=\gamma^{\prime}$. For the relevant $c$-factors, the distances to the end of the $C$-block are $d(\texttt{c4})=2m+k\text{ and }d(\texttt{c8})=m+\beta^{\prime},$ with corresponding lengths $|\texttt{c4}|=k-\beta^{\prime}\text{ and }|\texttt{c8}|=m.$
Throughout, “factor” refers to a symbolic word of the form $x^{y}$, while “factor set” refers to the induced set of game positions when a reference point is fixed. We only pass to factor sets when verifying reachability or anti-collision. To read the tables: fix a column factor set. Each row lists all potential factor-options of a relevant game value. Reachability tables ensure the existence of smaller values as move options, while anti-collision tables exclude equal values in the set of move options; here a triple of inequalities corresponds to the moves $s_{1},s_{2},s_{3}$, respectively.
Example A3 in bZB. This is the case where $\beta(i-1)=\alpha(i)=0$, and hence $\zeta(i-1)=0^{m}$. Moreover, it follows that $\alpha(i-1)=m-k$. The factor set A3 has size $m$ in such $B(i)$, and, by (14), we study the word
$B(i-1)\zeta(i-1)B(i)=1^{k}0^{m-k}2^{k}1^{m}3^{0}2^{m}0^{k}3^{m}\underline{0}^{m}1^{m}0^{0}\boldsymbol{2}^{m}1^{k}3^{m-k}2^{k}0^{m}3^{k},$ (8)
where the underlined 0-factor is $\zeta(i-1)$. Let us verify the factor set A3, induced by the bold font factor $\boldsymbol{2}^{m}$, assuming by induction that the previous factors are correctly defined and satisfy the mex-rule. Specifically, we have to prove that, for all positions in A3, the value is $v(2^{m})=2$.
By inspection, and since $\texttt{A2}=\emptyset$, all the $m$-options have value one, by induction. Note that in the case of sink preceding $B(0)$ (denoted sinkB), the related justification of A3 is immediate, since the move $m$ assures the value one, and both larger moves drown in the sink. So, let us focus on the “wrap case” of $B(i-1)\zeta(i-1)B(i)$, also denoted bZB; see Tables 2 and 8 for general reachability and anti-collision justifications, respectively.
What differs from the sink argument is that both long options have to be scrutinized for reachability (to find value 0) and for anti-collision (to avoid value 2).
Let us denote by $\langle x\rangle=\{1,\ldots,x\}$. The claim from Table 2 (reachability) is that both long factor-options contribute to the value zero. In fact, Z alone cannot compensate for the loss of some sink options. Instead, we use Table 2 to compute the ‘inverse images’ of both 0-factor-options Z and b4, and prove that their union is a superset of A3. We may take as reference point the first zero-value in b4 as position ‘one’. The 0-factor set b4 has size $m-\alpha^{\prime}=m-(m-k)=k$, while b5 has size $\gamma=[km]^{+}=m$ and the 0-factor set Z, which starts at position $m+k$, has size $m$. Thus, the candidate super-set is
$\Gamma$ $=(s_{3}+\langle k\rangle)\cup(m+k+s_{2}+\langle m\rangle)$
$=(3m+k+\langle k\rangle)\cup(m+k+2m+k+\langle m\rangle)$
$=\{3m+k+1,\ldots,3m+2k,3m+2k+1,\ldots,4m+2k\}.$
By using the same reference point, the positions of A3 are $\{3m+k+1,\ldots,4m+k\}\subset\Gamma$; namely, since $|\texttt{A1}|=m-\alpha=m$, the position just before A3 is $k+m+m+m=3m+k$. We may conclude that all positions in A3 have both values 0s and 1s as options.
Next we verify that, as claimed by Table 8 (anti-collision), the intersection of move-translated 2-factor positions with A3 is empty. There is exactly one 2-factor between a3 and A3, namely b3. The number of letters between b3 and A3 is
$|\texttt{b4}|+|\texttt{b5}|+|\texttt{Z}|+|\texttt{A1}|+|\texttt{A2}|$ $=m-\alpha(m-1)+\gamma(m-1)+\boldsymbol{m}+m-\alpha(0)+m-\gamma(0)$ (9)
$=k+m+m+m+m-k$ (10)
$=4m$ (11)
$>3m+k$ (12)
$=\max(S).$ (13)
Here the bold $m$ counts the number of symbols from the inserted $\zeta(i-1)=0^{m}$ factor. A consequence is that a3 is also sufficiently separated. By induction, this proves the correctness of the mex-values for the factor A3 in case bZB.
Example A3 in bB. This is the general case of consecutive $B$-blocks; see Table 3. We study the 2-factor set A3 in the word B(i1)B(i)B(i-1)B(i),
$1^{\gamma^{\prime}}3^{m-\gamma^{\prime}}2^{\gamma^{\prime}}0^{m-\alpha^{\prime}}3^{\gamma^{\prime}}\underline{1}^{m-\alpha}0^{\alpha}\boldsymbol{2}^{m-\alpha}1^{\gamma}3^{m-\gamma}2^{\gamma}0^{m-\alpha}3^{\gamma},$ (14)
where we have omitted the prefix A(i1)A(i-1), since it is separated by a block from A3. Similar to the bZB case, the move $s_{1}$ provides all $m-\alpha$ 11-values (note that in this case A2 is non-empty). But here, only one move suffices to reach the $m-\alpha$ 0-values, namely $s_{2}$. The candidate superset is $\Gamma=s_{2}+\langle m-\alpha^{\prime}\rangle$. By using the same reference position, the set A3 is:
$m-\alpha^{\prime}+\gamma^{\prime}+m-\alpha+\alpha+\langle m-\alpha\rangle$ $=2m+k+\langle m-\alpha\rangle$ (15)
$\subset\Gamma,$ (16)
by Lemma 5, and since $\alpha>\alpha^{\prime}$ (recall this is a case where $\beta(i-1)=\alpha(i)>0$). For anti-collision, by (14), the only $2$-factor set we have to test is b3, of size $\gamma^{\prime}$; see Table 9. We show that b3 fits between the moves $s_{2}$ and $s_{3}$. By Lemma 5, the number of letters between b3 and A3 is $2m+k$, which proves the first part. But $\gamma^{\prime}=\alpha$, so $\gamma^{\prime}+2m+k+m-\alpha=3m+k$, which implies that A3 does not collide with b3 via the move $s_{3}$.
Observe that this example resolves also A3 in bC, since $A(i)$ is a prefix in both the $B$ and $C$ blocks.
Example A3 in cC. Since $A(i)$ is a prefix in both the $B$ and $C$ blocks, this example resolves both A3 in cB and A3 in cC. Curiously, there are double entries for reachability in Table 6 (and the same in Table 4). The claimed moves for reachability to the two 0-factor sets c8 and c4 are $s_{2}$ and $s_{3}$, respectively. There are $m+\beta^{\prime}$ letters between c8 and A3. The candidate superset is
$\Gamma=(s_{3}+\langle k-\beta^{\prime}\rangle)\cup(m+k-\beta^{\prime}+s_{2}+\langle m\rangle),$ (17)
with $\beta^{\prime}<k$, since we study a $C$-block. Moreover, since $\alpha=\beta^{\prime}$, we get the interval
$\Gamma$ $=\{3m+k+1,\ldots,3m+k+k-\alpha\}\cup\{3m+2k-\alpha+1,\ldots,4m+2k-\alpha\}$ (18)
$=\{3m+k+1,\ldots,4m+2k-\alpha\}.$ (19)
And, with the same reference point, by inspection,
$2m+\beta^{\prime}+m+k-\beta^{\prime}+\texttt{A3}$ $=3m+k+\langle m-\alpha\rangle$ (20)
$=\{3m+k+1,\ldots,4m+k-\alpha\}\subset\Gamma.$ (21)
Thus reachability holds, and note that both moves $s_{3}$ and $s_{2}$ are required to 0-cover the factor.
Next, for anti-collision, as displayed in Table 12, the relevant 2-factors are c2 and c6. But, by the table, c2 is distanced by more than $3m+k$, while c6 is distanced exactly by $2m+k$, and has size $m-\alpha<m$, and hence neither $s_{3}$ collides with A3 in cC.
Example C9. In this example, we guide the reader through table justification. Note that C9 is distanced by more than $4m$ to its $A$-prefix. Hence, both reachability and anti-collision will be resolved locally within the sets C1 to C9. Thus, here we resolve this entry simultaneously in all four tables for bC and cC. Indeed, all Tables 5, 6, 11 and 12 have the same C9 entries. For reachability, $s1$ maps C9 into C8, $s2$ maps C9 onto C3 and $s3$ maps C9 into C2. The relevant sizes and distances are displayed in those tables, so verification is direct. Thus C9 satisfies reachability.
For anti-collision, both C5 and C7 are 3-factor sets, and require tests; see the last column and last two rows in Table 5. Observe that the size of $|\texttt{C8}|=m$, so $s_{1}$ does not collide with C7 and hence also not with C5. For $s_{2}$, note that $|\texttt{C5}\cup\texttt{C6}\cup\texttt{C7}\cup\texttt{C8}|=2m$, and hence the two last inequalities are also correct in both cells (since $k\geq 1$). Thus C9 does not collide.
Table 1: Reachability sinkB. This is the base case, sinkB, where the block $B(0)$ meets the sink. We omit the A2 column, since $\alpha(0)=0$. For each column factor, the table lists moves and target factors whose inverse images jointly cover the entire factor, ensuring reachability of the indicated value.
sinkB $\underset{m}{\texttt{A1}\text{``1''}}$ $\underset{m}{\texttt{A3}\text{``2''}}$ $\underset{k}{\texttt{B1}\text{``1''}}$ $\underset{m-k}{\texttt{B2}\text{``3''}}$ $\underset{k}{\texttt{B3}\text{``2''}}$ $\underset{m}{\texttt{B4}\text{``0''}}$ $\underset{k}{\texttt{B5}\text{``3''}}$
0
$s_{1}$
sink
$s_{2}$
sink
$s_{2}$
sink
$s_{3}$
sink
$s_{3}$
sink
$s_{1}$
B4
11
$s_{1}$
A1
$s_{2}$
A1
$s_{1}$, $s_{2}$
B1, A1
$s_{2}$
B1
22
$s_{1}$
A3
$s_{3}$
A3
Table 2: Reachability bZB. This is the wrap case, where the block $B(i)$ meets the inserted 0-factor $\zeta(i-1)=0^{m}$ preceded by B(i1)B(i-1). We omit the A2 column, since $\alpha(0)=0$. For each column factor, the table lists moves and target factors whose inverse images jointly cover the entire factor, ensuring reachability of the indicated value. For the factor b4, the distance to the current $B$-block is $d(\texttt{b4})=\gamma(m-1)=m$, and its length is $|\texttt{b4}|=m-\alpha(m-1)=k$.
bZB $\underset{m}{\texttt{A1}\text{``1''}}$ $\underset{m}{\texttt{A3}\text{``2''}}$ $\underset{k}{\texttt{B1}\text{``1''}}$ $\underset{m-k}{\texttt{B2}\text{``3''}}$ $\underset{k}{\texttt{B3}\text{``2''}}$ $\underset{m}{\texttt{B4}\text{``0''}}$ $\underset{k}{\texttt{B5}\text{``3''}}$
0
$s_{1}$
Z
$s_{2}$, $s_{3}$
Z, b4
$s_{2}$
Z
$s_{3}$
Z
$s_{3}$
Z
$s_{1}$
B4
11
$s_{1}$
A1
$s_{2}$
A1
$s_{1}$, $s_{2}$
B1, A1
$s_{2}$
B1
22
$s_{1}$
A3
$s_{3}$
A3
Table 3: Reachability bB. This is the generic case where the block $B(i)$ meets B(i1)B(i-1). For the factor b4, the distance to the current $B$-block is $d(\texttt{b4})=\gamma^{\prime}$, and its length is $|\texttt{b4}|=m-\alpha^{\prime}$.
bB $\underset{m-\alpha}{\texttt{A1}\text{``1''}}$ $\underset{\alpha}{\texttt{A2}\text{``0''}}$ $\underset{m-\alpha}{\texttt{A3}\text{``2''}}$ $\underset{\gamma}{\texttt{B1}\text{``1''}}$ $\underset{m-\gamma}{\texttt{B2}\text{``3''}}$ $\underset{\gamma}{\texttt{B3}\text{``2''}}$ $\underset{m-\alpha}{\texttt{B4}\text{``0''}}$ $\underset{\gamma}{\texttt{B5}\text{``3''}}$
0
$s_{1}$
b4
$s_{2}$
b4
$s_{1}$, $s_{2}$
A2, b4
$s_{3}$
b4
$s_{2}$, $s_{3}$
A2, b4
$s_{1}$, $s_{3}$
B4, A2
11
$s_{1}$
A1
$s_{2}$
A1
$s_{1}$, $s_{2}$
B1, A1
$s_{2}$
B1
22
$s_{1}$
A3
$s_{1}$, $s_{3}$
B3, A3
Table 4: Reachability cB. This is the generic case where the block $B(i)$ meets C(i1)C(i-1), which implies $\beta(i-1)<\alpha(i-1)$. For the relevant $c$-factors, the distances to the end of the $C$-block are $d(\texttt{c4})=2m+\beta^{\prime}$ and $d(\texttt{c8})=\beta^{\prime},$ with corresponding lengths $|\texttt{c4}|=k-\beta^{\prime}$ and $|\texttt{c8}|=m.$
cB $\underset{m-\alpha}{\texttt{A1}\text{``1''}}$ $\underset{\alpha}{\texttt{A2}\text{``0''}}$ $\underset{m-\alpha}{\texttt{A3}\text{``2''}}$ $\underset{\gamma}{\texttt{B1}\text{``1''}}$ $\underset{m-\gamma}{\texttt{B2}\text{``3''}}$ $\underset{\gamma}{\texttt{B3}\text{``2''}}$ $\underset{m-\alpha}{\texttt{B4}\text{``0''}}$ $\underset{\gamma}{\texttt{B5}\text{``3''}}$
0
$s_{1}$, $s_{2}$
c8, c4
$s_{2}$, $s_{3}$
c8, c4
$s_{1}$, $s_{2}$
A2, c8
$s_{3}$
c8
$s_{2}$, $s_{3}$
A2, c8
$s_{1}$, $s_{3}$
B4, A2
11
$s_{1}$
A1
$s_{2}$
A1
$s_{1}$, $s_{2}$
B1, A1
$s_{2}$
B1
22
$s_{1}$
A3
$s_{1}$, $s_{3}$
B3, A3
Table 5: Reachability bC. This is the generic case, where the block $C(i)$ meets B(i1)B(i-1), which implies $\beta(i)<\alpha(i)$. For the factor b4, the distance to the current $B$-block is $d(\texttt{b4})=\gamma^{\prime}$, and its length is $|\texttt{b4}|=m-\alpha^{\prime}$.
bC $\underset{m-\alpha}{\texttt{A1}\text{``1''}}$ $\underset{\alpha}{\texttt{A2}\text{``0''}}$ $\underset{m-\alpha}{\texttt{A3}\text{``2''}}$ $\underset{m}{\texttt{C1}\text{``1''}}$ $\underset{m}{\texttt{C2}\text{``2''}}$ $\underset{\beta}{\texttt{C3}\text{``1''}}$ $\underset{k-\beta}{\texttt{C4}\text{``0''}}$ $\underset{m-k}{\texttt{C5}\text{``3''}}$ $\underset{\beta}{\texttt{C6}\text{``2''}}$ $\underset{k-\beta}{\texttt{C7}\text{``3''}}$ $\underset{m}{\texttt{C8}\text{``0''}}$ $\underset{\beta}{\texttt{C9}\text{``3''}}$
0
$s_{1}$
b4
$s_{2}$
b4
$s_{1}$, $s_{2}$
A2, b4
$s_{2}$, $s_{3}$
A2, b4
$s_{2}$
A2
$s_{3}$
A2
$s_{3}$
A2
$s_{1}$
C4
$s_{1}$
C8
11
$s_{1}$
A1
$s_{1}$, $s_{2}$
C1, A1
$s_{2}$
C1
$s_{1}$, $s_{2}$
C3, C1
$s_{2}$
C1
$s_{2}$
C3
22
$s_{1}$
C2
$s_{3}$
A3
$s_{3}$
C2
Table 6: Reachability cC. This is the generic case, where the block $C(i)$ meets C(i1)C(i-1), which implies $\beta(i)<\alpha(i)$ and $\beta(i-1)<\alpha(i-1)$. For the relevant $c$-factors, the distances to the end of the $C$-block are $d(\texttt{c4})=2m+\beta^{\prime}$ and $d(\texttt{c8})=\beta^{\prime},$ with corresponding lengths $|\texttt{c4}|=k-\beta^{\prime}$ and $|\texttt{c8}|=m.$
cC $\underset{m-\alpha}{\texttt{A1}\text{``1''}}$ $\underset{\alpha}{\texttt{A2}\text{``0''}}$ $\underset{m-\alpha}{\texttt{A3}\text{``2''}}$ $\underset{m}{\texttt{C1}\text{``1''}}$ $\underset{m}{\texttt{C2}\text{``2''}}$ $\underset{\beta}{\texttt{C3}\text{``1''}}$ $\underset{k-\beta}{\texttt{C4}\text{``0''}}$ $\underset{m-k}{\texttt{C5}\text{``3''}}$ $\underset{\beta}{\texttt{C6}\text{``2''}}$ $\underset{k-\beta}{\texttt{C7}\text{``3''}}$ $\underset{m}{\texttt{C8}\text{``0''}}$ $\underset{\beta}{\texttt{C9}\text{``3''}}$
0
$s_{1}$, $s_{2}$
c8, c4
$s_{2}$, $s_{3}$
c8, c4
$s_{1}$, $s_{2}$
A2, c8
$s_{2}$, $s_{3}$
A2, c8
$s_{2}$
A2
$s_{3}$
A2
$s_{3}$
A2
$s_{1}$
C4
$s_{1}$
C8
11
$s_{1}$
A1
$s_{1}$, $s_{2}$
C1, A1
$s_{2}$
C1
$s_{1}$, $s_{2}$
C3, C1
$s_{2}$
C1
$s_{2}$
C3
22
$s_{1}$
C2
$s_{3}$
A3
$s_{3}$
C2
Table 7: Anti-collision sink B. This table concerns the first $B$-block sink B. Each cell lists a triple of inequalities corresponding to the moves $s_{1}$, $s_{2}$, and $s_{3}$, respectively. For example (with a fixed reference point), $\max(\texttt{A1})+s_{1}<\min(\texttt{B1})$, whereas $\min(\texttt{A1})+s_{>}\max(\texttt{B1})$. Subscripts indicate the lengths of the corresponding factors. In each cell, the first occurrence of the inequality symbol “>>” is tight.
sink B $\underset{m}{\texttt{A1}\text{``1''}}$ $\underset{0}{\texttt{A2}\text{``0''}}$ $\underset{m}{\texttt{A3}\text{``2''}}$ $\underset{k}{\texttt{B1}\text{``1''}}$ $\underset{m-k}{\texttt{B2}\text{``3''}}$ $\underset{k}{\texttt{B3}\text{``2''}}$ $\underset{m}{\texttt{B4}\text{``0''}}$ $\underset{k}{\texttt{B5}\text{``3''}}$
value 0
[…]
value 1
[…]
value 2
[…]
value 3
[…]
Table 8: Anti-collision bZB. This table concerns the block concatenation $B(i-1)\zeta(i-1)B(i)$, whenever $\alpha(i)=0$. Thus, Z denotes the set of positions induced by the factor $\zeta(i-1)=0^{m}$. Some row factors have distance greater than $3m+k$ from the factor under consideration in the column. In such cases, any preceding row factor with the same game value is omitted. For example, $\min(\texttt{Z}+3m+k)>\max(\texttt{B4})$, and hence the factor b4 of the same value may be omitted. Recall that factors belonging to the current block are denoted by capital letters, while factors from the preceding block are denoted by lowercase letters. Row headers list the factor name and, for factors in the preceding block, the distance to the end of the block, together with the factor length; for factors in the current block, only the factor length is displayed. Note that the factors b2 and A2 are empty.
bZB $\underset{m}{\texttt{Z}\text{``0''}}$ $\underset{m}{\texttt{A1}\text{``1''}}$ $\underset{0}{\texttt{A2}}$ $\underset{m}{\texttt{A3}\text{``2''}}$ $\underset{k}{\texttt{B1}\text{``1''}}$ $\underset{m-k}{\texttt{B2}\text{``3''}}$ $\underset{k}{\texttt{B3}\text{``2''}}$ $\underset{m}{\texttt{B4}\text{``0''}}$ $\underset{k}{\texttt{B5}\text{``3''}}$
value 0
[…]
value 1
[…]
value 2
[…]
value 3
[…]
Table 9: Anti-collision bB. This table concerns the general block concatenation bB. The factor sets a1, a2, a3 and b2 are omitted since they are of at least a block length to the first possible collision, respectively.
bB $\underset{m-\alpha}{\texttt{A1}\text{``1''}}$ $\underset{\alpha}{\texttt{A2}\text{``0''}}$ $\underset{m-\alpha}{\texttt{A3}\text{``2''}}$ $\underset{\gamma}{\texttt{B1}\text{``1''}}$ $\underset{m-\gamma}{\texttt{B2}\text{``3''}}$ $\underset{\gamma}{\texttt{B3}\text{``2''}}$ $\underset{m-\alpha}{\texttt{B4}\text{``0''}}$ $\underset{\gamma}{\texttt{B5}\text{``3''}}$
value 0
[…]
value 1
[…]
value 2
[…]
value 3
[…]
Table 10: Anti-collision cB. This table concerns the block concatenation cB. The factors of the $B$-block are omitted since they were already tested in Table 9.
cB $\underset{m-\alpha}{\texttt{A1}\text{``1''}}$ $\underset{\alpha}{\texttt{A2}\text{``0''}}$ $\underset{m-\alpha}{\texttt{A3}\text{``2''}}$ $\underset{\gamma}{\texttt{B1}\text{``1''}}$ $\underset{m-\gamma}{\texttt{B2}\text{``3''}}$ $\underset{\gamma}{\texttt{B3}\text{``2''}}$ $\underset{m-\alpha}{\texttt{B4}\text{``0''}}$ $\underset{\gamma}{\texttt{B5}\text{``3''}}$
value 0
[…]
value 1
[…]
value 2
[…]
value 3
[…]
Table 11: Anti-collision bC. This table concerns the block concatenation bC. The factor sets a1, a2 and a3 have been omitted since they are a block length apart their corresponding values in the $C$-block.
bC $\underset{m-\alpha}{\texttt{A1}\text{``1''}}$ $\underset{\alpha}{\texttt{A2}\text{``0''}}$ $\underset{m-\alpha}{\texttt{A3}\text{``2''}}$ $\underset{m}{\texttt{C1}\text{``1''}}$ $\underset{m}{\texttt{C2}\text{``2''}}$ $\underset{\beta}{\texttt{C3}\text{``1''}}$ $\underset{k-\beta}{\texttt{C4}\text{``0''}}$ $\underset{m-k}{\texttt{C5}\text{``3''}}$ $\underset{\beta}{\texttt{C6}\text{``2''}}$ $\underset{k-\beta}{\texttt{C7}\text{``3''}}$ $\underset{m}{\texttt{C8}\text{``0''}}$ $\underset{\beta}{\texttt{C9}\text{``3''}}$
value 0
[…]
value 1
[…]
value 2
[…]
value 3
[…]
Table 12: Anti-collision cC. This table concerns the block concatenation cC. The factors from the current $C$-factor are omitted since they were already tested in Table 11.
cC $\underset{m-\alpha}{\texttt{A1}\text{``1''}}$ $\underset{\alpha}{\texttt{A2}\text{``0''}}$ $\underset{m-\alpha}{\texttt{A3}\text{``2''}}$ $\underset{m}{\texttt{C1}\text{``1''}}$ $\underset{m}{\texttt{C2}\text{``2''}}$ $\underset{\beta}{\texttt{C3}\text{``1''}}$ $\underset{k-\beta}{\texttt{C4}\text{``0''}}$ $\underset{m-k}{\texttt{C5}\text{``3''}}$ $\underset{\beta}{\texttt{C6}\text{``2''}}$ $\underset{k-\beta}{\texttt{C7}\text{``3''}}$ $\underset{m}{\texttt{C8}\text{``0''}}$ $\underset{\beta}{\texttt{C9}\text{``3''}}$
value 0
[…]
value 1
[…]
value 2
[…]
value 3
[…]
It remains to justify the general case of item (ii), for a fixed $k$, and when $\delta=2nm+k$, for $n\geq 1$. For each increase in $n$, the period length increases by $4m^{2}/{\gcd(m,\delta)}$, and this is due to the fact that each block is increased by exactly $4m$ positions via the two concatenated factors $0^{m}1^{m}$ and $2^{m}3^{m}$; see Figure 4. The justification is similar to case (i), but here it invokes another quadratic term. Apart from this observation, the mex verification remains the same as for the case $n=1$. This concludes the proof of Theorem 3.
Figure 4: Nimbers for $m=3$, $\delta\equiv 4\pmod{6}$ (left) and $m=4$ with $\delta\equiv 7\pmod{8}$ (right).
3 Final remarks
On a more philosophical matter, with respect to the ease of the proof of Theorem 2, one might ponder the value of understanding the fine details of a small sub-family of subtraction games. However this type of result has been on the list of conjectures for a long time, and our conjectured reciprocity with the wall convention requires further research. For example, Richard Guy writes in the first problem, in the first edition of Unsolved problems in combinatorial games [4] “Subtraction games are known to be periodic. Investigate the relationship between the subtraction set and the length and structure of the period.”; see also [6]. The second part of the conjecture for three move games [2, 7] concerns the non-additive three-move rulesets. Although they all are conjectured to have period length “a sum of two moves”, and any such instance would individually be easily verified, the full classification appears to satisfy some fractal structure [2], and has not yet been fully understood. See also the recent survey on subtraction games [5].
Acknowledgements. Thanks to ChatGPT 5.2 for the two images in Figure 1.

### v2

@@
 Subtraction games are a classical topic in Combinatorial Game Theory.
-A result of Golomb (1966) shows that every subtraction game with a finite move set has an eventually periodic nim-sequence, but the known proof yields only an exponential upper bound on the period length.
-Flammenkamp (1996) conjectures a striking classification for three-move subtraction games: non-additive rulesets exhibit linear period lengths of the form “the sum of two moves”, where the choice of which two moves displays fractal-like behavior, while additive sets $S=\{a,b,a+b\}$ have purely periodic outcomes with linear or quadratic period lengths.
+An early result of Golomb (1966) shows that every subtraction game with a finite move set has an eventually periodic nim-sequence, but the known proof yields only an exponential upper bound on the period length.
+Even three-move games are not yet fully understood.
+Nevertheless, Flammenkamp (1997) conjectures a striking classification: non-additive rulesets exhibit linear period lengths of the form “the sum of two moves”, where the choice of which two moves displays fractal-like behavior, while additive sets $S=\{a,b,a+b\}$ have purely periodic outcomes with linear or quadratic period lengths.
 Despite early attention in Winning Ways (1982), the general additive case remains open.
-We introduce and analyze a dual winning convention, which we call sink subtraction.
+We analyze a dual winning convention, called sink subtraction.
 Unlike the standard wall convention, where moves to negative positions are forbidden, the sink convention declares a player the winner upon moving to a non-positive position.
-We show that additive sink subtraction admits a complete solution: the nim-sequence is purely periodic with an explicit linear or quadratic period formula, and we conjecture a duality between additive sink subtraction and classical wall subtraction.
+We show that additive sink subtraction admits a complete solution: its nim-sequence is purely periodic with an explicit linear or quadratic period formula.
+We further conjecture a nimber duality between additive sink subtraction and classical wall subtraction.
 Keywords: additive subtraction game, nimber, periodicity, sink convention.
@@
 For any $x\in\mathbb{N}$ and all $s\in S$, $x-s\in\mathbb{Z}$ is a move option.
-The sink is an infinite sequence of terminal P-positions for all non-positive integers.
+The sink is an infinite sequence of terminal $\mathcal{P}$-positions ($\mathcal{P}$revious player winning positions) for all non-positive integers.
 No move is possible from a position in the sink.
@@
 To the Right, the boy has only one move, because the larger move is prohibited by the wall.
-Thus the wall-game is determined already by parity.
+Here, the wall game is determined already by parity.
 Remark 1.
@@
 nim-value) is defined via the standard minimum exclusive rule; for any finite $X\subset\mathbb{N}_{0}$, let $\mathrm{mex}X=\min\{\mathbb{N}_{0}\setminus X\}$, where $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$.
-Then the nim value algorithm of $x$ is $v(x)=\mathrm{mex}\{v(x^{\prime})\}$, where $x^{\prime}$ ranges over the options of $x$.
+Then the nim-value algorithm of $x$ is $v(x)=\mathrm{mex}\{v(x^{\prime})\}$, where $x^{\prime}$ ranges over the options of $x$.
 Note that this gives value zero for all positions in the sink.
@@
 $1,1,2,{\bf 2,1,0,0,1,1,0},2,1,0,0,1,1,\ldots$
-with a pre-period of size three but the same period length 7.
-The periodic patterns are the same but occur in a different order: 0011021 and 2100110 respectively.
-To the best of our knowledge, sink subtraction has not yet been studied in the literature.
-It is the purpose of this note to enter this unpaved territory.
-We use the convention of $x=0$ belonging to the sink, so we start nim-sequences by the (non-zero) value of position one.
+with a pre-period of size three but the same period length 7. The periodic patterns are the same but occur in a different order: 0011021 and 2100110 respectively.
+To the best of our knowledge, sink subtraction has only been studied in the literature in terms of outcomes, and only in one paper [1].
+They remark that one can relate the outcomes of the two conventions via the formula
+$o_{\rm wall}(x)=o_{\rm sink}(x+\max(S)+1).$ (1)
+[…]
+In this paper we continue the pursuit of [1], by studying the more general nim-values of sink subtraction.
+We use the convention of position $x=0$ belonging to the sink, so we start nim-sequences by the value (one) of position one.
 Of course, the folklore periodicity theorem holds for rulesets where the terminal condition is a sink instead of a ‘wall’, with exactly the same proof.
@@
 There is a recent paper that discusses such settings [miklós2023superpolynomial], and they establish super-polynomial period length for three-move games with specific such relaxed winning conditions.
-The purpose of this note is to establish the period lengths of the ruleset family additive sink subtraction $S$ with subtraction set
+The purpose of this paper is to establish the period lengths of the ruleset family additive sink subtraction $S$ with subtraction set
 $S(m,\delta)=\{m,m+\delta,2m+\delta\},$ (2)
 where $m,\delta\in\mathbb{N}$.
-There is a conjecture from Flammenkamp’s PhD thesis [2] concerning the periodicity of Additive Wall Subtraction, and this is to some degree paraphrasing the treatment in Winning Ways [1], where the linear-period-length part of the conjecture has been proved.
+There is a conjecture from Flammenkamp’s PhD thesis [3] concerning the periodicity of additive wall subtraction, and this partially paraphrases the treatment in Winning Ways [2], where the linear-period-length part of the conjecture has been proved.
 As far as the authors know, the non-linear (quadratic) part has not yet been established.
@@
 Consider additive sink subtraction with subtraction set $S(m,\delta)=\{m,m+\delta,2m+\delta\}$, and let $d:=\delta\ \pmod{2m}$, with $0\leq d<2m$.
-Then $S$ is purely periodic, and its period length is
+Then $S$ has purely periodic nim-values, with period length
 $p(m,\delta)=\begin{cases}3m+2\delta-d,&\mathrm{if}\ d\in[0,m];\\ m(m+2\delta+d)/\gcd(m,d),&\mathrm{otherwise}.\end{cases}$
+The immediate consequence for the outcomes of additive wall subtraction is as follows.
+Corollary 4.
+Let $m,\delta\in\mathbb{N}$, and consider additive wall subtraction with subtraction set $S(m,\delta)=\{m,m+\delta,2m+\delta\}$.
+Then the period length of the outcomes is the same as in Theorem 3, $p(m,\delta)$.
+Proof.
+This is obvious by (1). ∎
+We conjecture that there is an explicit duality formula for relating the nimbers of additive sink subtraction with those of additive wall subtraction, of the form, for all $x$,
+$v_{\rm wall}\bigl([x]_{p}\bigr)=\sigma\bigl(v_{\rm sink}\bigl([s_{2}-x]_{p}\bigr)\bigr),$ (3)
+where $\sigma:0\rightarrow 2,1\rightarrow 1,2\rightarrow 0,3\rightarrow 3$. Here $x]_{p}$ denotes the heap size modulo the (outcome) period length $p$.
+Example 5.
+The wall and sink nim-sequences of $S=\{2,5,7\}$ are displayed in the below table; they have period length $22$.
+The upper (lower) row represents the heap sizes for wall (sink), and the middle row entries illustrate the respective nimbers (wall upper and sink lower).
+Note that the values satisfy Equation 3.
+$x_{\mathrm{wall}}$ 0 11 22 33 44 55 66 77 88 99 $10$ $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$ 1919 2020 2121
+$v(x)$ $0\atop 2$ $0\atop 2$ 11 11 $0\atop 2$ $2\atop 0$ 11 33 $2\atop 0$ $2\atop 0$ $0\atop 2$ 33 11 $0\atop 2$ $0\atop 2$ 11 11 $2\atop 0$ $2\atop 0$ 33 33 $2\atop 0$
+$x_{\mathrm{sink}}$ $17$ $16$ $15$ $14$ $13$ $12$ $11$ $10$ 99 88 77 66 55 44 33 22 11 $22$ 2121 2020 1919 $18$
 2 A proof of the main result
@@
 For convenience we restrict $m<\delta<2m$, and let $\delta=m+k$.
-(The general case turns out to be an immediate consequence.) As an example, let us begin by describing the periodic word for any rulesets of the form “the lower case”, i.e.
+(The general case turns out to be an immediate consequence.) Similar to case (i), let the alphabet be $\{0,1,2,3\}$.
+As an example, let us begin by describing the periodic word for any rulesets of the form “the lower case”, i.e.
 whenever $k=1$.
@@
 for $0\leq i\leq m-1$, and at last, we concatenate the factor $0^{m}$.
+[^2] The period length for this case is mentioned in [1], but without giving the explicit block construction or a proof.
 By increasing the ‘block speed’, the factor $1\leq k<m$, another type of blocks appears, the “$C$-blocks”, and the extreme “upper case” will favor only the $C$-blocks.
 The challenge is to understand the intermediate cases, with mixing between the block types; in Figure 2, we illustrate all periods for the case of $m=19$, $0<k<19$.
-Figure 2: Nimber periods for $m=19$ and $0<k<19$, starting from $x=1$. The colors correspond to the nimbers; purple (0), blue (1), green (2) and yellow (3). Zoom in for details.
+Figure 2: Nimber periods for $m=19$ and $0<k<19$, starting from $x=1$.
+The colors correspond to the nimbers; purple (0), blue (1), green (2) and yellow (3).
+The rows have different scaling as to fit the full periods in each case.
+Zoom in for details.
 Let $x]_{y}$ denote the smallest nonnegative representative of $x\pmod{y}$.
@@
 That is, if $W=(w_{i})_{i\geq 1}$ is the infinite word on the same alphabet such that, for all $i\geq 1$, $w_{i}$ is the nimber of $S$, then $W=PP\ldots$.
-Observation 4.
+Observation 6.
 In spite of the periodicity hinging on a quadratic expression in $m$, we are able to verify correctness in constant time.
@@
 If $m=6$ and $k=4$, then the block structure is $BCB0^{m}BCB0^{m}$, with claimed period length $2(4\cdot 6+2c-2a)+6\cdot 6+4+b-a+6=108$.
-The cases for $m=6$ are illustrated in the lower picture of Figure 3.
+The cases for $m=6$ are illustrated in the middle picture of Figure 3.
 Recall that we here assume $\delta=m+k$, $1\leq k<m$.
@@
 $=\frac{m(4m+3k)}{\gcd(m,k)}.$
-The blocks use two independent indices; the second one will be used throughout the tables to come.
+We use elementary algebra to justify mex-consistency.
+Each maximum factor $F$, of $P$, is of the form $F=a^{b}$ with $a\in\{0,1,2,3\}$ and where $0\leq b\leq m$ is maximum such that adjacent letters differ from $a$; the length of $F$ is $|F|=b$.
+We often omit the word ‘maximum’ and call a maximum factor simply by a factor.
+The factor-value of a factor $F$ is $v(F)=a$ if $F=a^{b}$.
+We will treat each factor as a single entity, and use terminology such as factor-option and factor-value.
+The advantage is that it drastically reduces the number of cases, but an apparent disadvantage is that sometimes individual positions within a factor can have options in distinct factors, for a given move.
+This phenomenon proves unavoidable though, and is part and parcel of the quadratic period length, as we will see.
+By intuition, when factor relations settle, then we approach the full period.
+The blocks will use two independent indices, the latter of which will enumerate its factors.
   1. 1.
@@
     Within a fixed block, we label its consecutive factors by a local index.
-For a $B$-block, the factors are denoted $\texttt{B1},\ldots,\texttt{B5}$, where Bj refers to the (j+3)(j+3)rd factor of that block, for $j\in\{1,\ldots,5\}$.
-Similarly, Aj denotes the $j$th factor of the prefix $A(i)$ associated with the same block index $i$.
+For a given $B(i)$-block, the factors are denoted $A_{1},A_{2},A_{3},B_{1},\ldots,B_{5}$, where $A_{j}$ denotes the $j$th factor of the prefix $A(i)$ and $B_{j}$ refers to the (j+3)(j+3)rd factor of that block.
 Unless explicitly stated otherwise, the block index $i$ is fixed when referring to factor indices.
-For example, the notation B3 always refers to the sixth factor within a single $B(i)$ block, independent of the block index $i$.
+For example, the notation $B_{3}$ always refers to the sixth factor within a single $B(i)$ block, where the block index $i$ is given by the surrounding context.
 Whenever we refer to an object other than a block or a factor, we explicitly use the term word.
 Recall that the block length of $B$ is $4m-2\alpha+2\gamma=4m+2k>3m+k$ (by $\gamma>\alpha$), while the block length of $C$ is $6m+k-\alpha+\beta=5m+2k>3m+k$ (by $\beta<\alpha$).
-One can check that, under block concatenation, the distance between two similar factors is always greater than $3m+k$; two A factors are similar, if they carry the same j index, and analogously for B and C.
-We will treat each factor as a single entity, and use terminology such as factor-option and factor value.
-The advantage is that it drastically reduces the number of cases, but an apparent disadvantage is that sometimes individual positions within a factor can have options in distinct factors, for a given move.
-This phenomenon proves unavoidable though, and is part and parcel of the quadratic period length, as we will see.
-By intuition, when factor relations settle, then we approach the full period.
-We use elementary algebra to justify mex-consistency.
-Each factor $F$, of $P$, is of the form $F=x^{y}$ with $x\in\{0,1,2,3\}$ and where $y\geq 0$ is maximum such that adjacent letters differ from $x$; the length of $F$ is $|F|=y$.
-The factor value of a factor $F$ is $v(F)=x$ if $F=x^{y}$.
-We often want to identify a factor with its corresponding set of positions.
-To this purpose it is convenient to temporarily identify position one with the first letter of some relevant factor, and we write F to denote the factor set, the set of such induced positions corresponding to a factor $F$.
-The value of a factor set is its factor value.
-A factor-option of $F$ is a factor $F^{\prime}$ such that there is an $x\in\texttt{F'}$ such that, for some $y\in\texttt{F}$, $x-y\in\{m,2m+k,3m+k\}$.
-Any such terminology is local and used as a computational device.
+One can check that, under block concatenation, the distance between two similar factors is always greater than the largest move, $3m+k$; two $A$-factors are similar, if they carry the same $j$-index, and analogously for $B$- and $C$-factors.
 We display tables with factor-options and their values.
@@
 Tables 7 to 12 verify anti-collision, the second mex-principle, that any conjectured factor-value, cannot have the same value at the level of any factor-option.
+We often study the corresponding set of positions of a given factor.
+To this purpose it is convenient to temporarily identify position one with the first letter of some relevant factor, and we use a special font F to denote the factor set, the set of such induced positions corresponding to a factor $F$.
+Any such notation is local and used as a computational device.
+Each factor is then interpreted as a set of consecutive positions, all carrying the same game value, determined relative to this reference point.
+In this way, the symbolic word representations, together with the specified factor lengths, are mapped unambiguously to actual game positions, allowing us to analyze move options between factors.
 Throughout the tables, the symbols $s_{1}$, $s_{2}$, and $s_{3}$ denote the moves $m$, $2m+k$, and $3m+k$, respectively.
@@
 In tables for concatenations (e.g.
-BB), we use uppercase factor names for the current block $B(i)$ and lowercase for the preceding block B(i1)B(i-1).
-For example, B4 is the set of positions corresponding to the factor $0^{m-\alpha}$ in $B(i)$, while b4 is the set of positions corresponding to the factor $0^{m-\alpha^{\prime}}$ in B(i1)B(i-1); Throughout, we write $\alpha^{\prime}=\alpha(i-1)$, $\beta^{\prime}=\beta(i-1)$ and $\gamma^{\prime}=\gamma(i-1)$.
-In the tables, rather than fixing a specific heap of size one, we adopt the following convention.
-For each row (i.e.
-each game value under consideration), we choose as reference point the smallest position in the leftmost factor of that value; for the sink we use position 0 as reference.
-Each factor is then interpreted as a set of consecutive positions, all carrying the same game value, determined relative to this reference point.
-In this way, the symbolic word representations, together with the specified factor lengths, are mapped unambiguously to actual game positions, allowing us to analyze move options between factors.
+bB concerning reachability in (16)), we use an uppercase letter for the current block $B(i)$ and a lowercase for the preceding block B(i1)B(i-1).
+For example, the factor set B4 is the set of positions corresponding to the factor $0^{m-\alpha}$ in $B(i)$, while b4 is the set of positions corresponding to the factor $0^{m-\alpha^{\prime}}$ in B(i1)B(i-1); throughout the tables, we write $\alpha^{\prime}=\alpha(i-1)$, $\beta^{\prime}=\beta(i-1)$ and $\gamma^{\prime}=\gamma(i-1)$.
 In the reachability tables, distances are measured from the beginning of the relevant factor to the end of the preceding block.
-For the factor b4, the distance to the end of the $B$-block is $d(\texttt{b4})=m-\gamma^{\prime}+\alpha^{\prime}=m-k$ (by Lemma 5), and its length is $|\texttt{b4}|=\gamma(i-1)=\gamma^{\prime}$.
-For the relevant $c$-factors, the distances to the end of the $C$-block are $d(\texttt{c4})=2m+k\text{ and }d(\texttt{c8})=m+\beta^{\prime},$ with corresponding lengths $|\texttt{c4}|=k-\beta^{\prime}\text{ and }|\texttt{c8}|=m.$
-Throughout, “factor” refers to a symbolic word of the form $x^{y}$, while “factor set” refers to the induced set of game positions when a reference point is fixed.
-We only pass to factor sets when verifying reachability or anti-collision.
+For example, the distance to the end of the $B$-block of the factor b4 is $d(\texttt{b4})=m-\gamma^{\prime}+\alpha^{\prime}=m-k$ (by Lemma 7), and its length is $|\texttt{b4}|=\gamma(i-1)=\gamma^{\prime}$.
+Similarly, $d(\texttt{c4})=2m+k\text{ and }d(\texttt{c8})=m+\beta^{\prime},$ with corresponding lengths $|\texttt{c4}|=k-\beta^{\prime}\text{ and }|\texttt{c8}|=m.$
+The value of a factor set is its factor-value.
+A factor-option of $F$ is a factor $F^{\prime}$ such that there is an $x\in\texttt{F'}$ such that, for some $y\in\texttt{F}$, $x-y\in\{m,2m+k,3m+k\}=\{s_{1},s_{2},s_{3}\}$.
 To read the tables: fix a column factor set.
@@
 Reachability tables ensure the existence of smaller values as move options, while anti-collision tables exclude equal values in the set of move options; here a triple of inequalities corresponds to the moves $s_{1},s_{2},s_{3}$, respectively.
-Example A3 in bZB.
+Example A3 in sinkB vs.
+bZB.
 This is the case where $\beta(i-1)=\alpha(i)=0$, and hence $\zeta(i-1)=0^{m}$.
@@
 3 Final remarks
-On a more philosophical matter, with respect to the ease of the proof of Theorem 2, one might ponder the value of understanding the fine details of a small sub-family of subtraction games.
+On a more philosophical matter, with respect to the simplicity of the proof of Theorem 2, one might ponder the value of understanding the fine details of a small sub-family of subtraction games.
 However this type of result has been on the list of conjectures for a long time, and our conjectured reciprocity with the wall convention requires further research.
@@
 See also the recent survey on subtraction games [6].
-Acknowledgements. Thanks to ChatGPT 5.2 for the two images in Figure 1.
+Acknowledgements. We thank Richard Nowakowski for pointing out the reference [1]. The images in Figure 1 were generated with assistance from ChatGPT 5.2.

### v3

@@
 Even three-move games are not yet fully understood.
-Nevertheless, Flammenkamp (1997) conjectures a striking classification: non-additive rulesets exhibit linear period lengths of the form “the sum of two moves”, where the choice of which two moves displays fractal-like behavior, while additive sets $S=\{a,b,a+b\}$ have purely periodic outcomes with linear or quadratic period lengths.
+Nevertheless, Flammenkamp (1997) conjectures a striking classification: non-additive rulesets have linear period lengths of the form “the sum of two moves”, where the choice of which two moves displays fractal-like behavior, while additive sets $S=\{a,b,a+b\}$ have purely periodic outcomes with linear or quadratic period lengths.
 Despite early attention in Winning Ways (1982), the general additive case remains open.
@@
 Note that this gives value zero for all positions in the sink.
-For example, if $S=\{2,5\}$, then starting from $x=0$, ‘wall’ has a purely periodic nim-sequence of period length 7,
+For example, if $S=\{2,5\}$, then starting from $x=0$, wall has a purely periodic nim-sequence of period length 7,
 ${\bf 0,0,1,1,0,2,1},0,0,1,1,0,2,1\ldots,$
-while ‘sink’, starting from $x=1$, begins
+while sink, starting from $x=1$, begins
 $1,1,2,{\bf 2,1,0,0,1,1,0},2,1,0,0,1,1,\ldots$
@@
 The purpose of this paper is to establish the period lengths of the ruleset family additive sink subtraction $S$ with subtraction set
-$S(m,\delta)=\{m,m+\delta,2m+\delta\},$ (2)
-where $m,\delta\in\mathbb{N}$.
-There is a conjecture from Flammenkamp’s PhD thesis [3] concerning the periodicity of additive wall subtraction, and this partially paraphrases the treatment in Winning Ways [2], where the linear-period-length part of the conjecture has been proved.
-As far as the authors know, the non-linear (quadratic) part has not yet been established.
+$S=\{s_{1},s_{2},s_{1}+s_{2}\},$ (2)
+where $s_{1},s_{2}\in\mathbb{N}$, $s_{1}<s_{2}$.
+There is a conjecture from Flammenkamp’s PhD thesis [3] concerning the periodicity of additive wall subtraction, and this partially paraphrases the treatment in Winning Ways [2], where the linear-period-length part of the conjecture has been explained (without a proof).
+As far as the authors know, the non-linear (quadratic) part has not yet been fully solved in the literature, but conjectured also in [9].
 [^1] Non-additive three-move games are conjectured to have period lengths of the form “the sum of two moves”, where “which two moves?” appear to satisfy some fractal patterns [3].
-Let us state our main result.
+Recently a subset of the authors completely determine the closed formulas of the nimbers in the primitive quadratic regime [6] (whenever $a<\delta<2a$ and $\gcd{a,\delta}=1$), by applying elementary combinatorial number theory to the $\mathcal{P}$-position bracket expression from [2] (see also Section 3).
+Let us state our main result. In the statement we use a different representation of (2).
 Theorem 3 (Main Theorem).
@@
 We conjecture that there is an explicit duality formula for relating the nimbers of additive sink subtraction with those of additive wall subtraction, of the form, for all $x$,
-$v_{\rm wall}\bigl([x]_{p}\bigr)=\sigma\bigl(v_{\rm sink}\bigl([s_{2}-x]_{p}\bigr)\bigr),$ (3)
+$v_{\rm wall}\bigl([x]_{p}\bigr)=\sigma\bigl(v_{\rm sink}\bigl([-s_{2}-x]_{p}\bigr)\bigr),$ (3)
 where $\sigma:0\rightarrow 2,1\rightarrow 1,2\rightarrow 0,3\rightarrow 3$. Here $x]_{p}$ denotes the heap size modulo the (outcome) period length $p$.
@@
 $x_{\mathrm{sink}}$ $17$ $16$ $15$ $14$ $13$ $12$ $11$ $10$ 99 88 77 66 55 44 33 22 11 $22$ 2121 2020 1919 $18$
+Let us remark that the well-known Ferguson pairing principle [2] fails to hold in sink subtraction.
+Remark 6.
+Let us first reiterate the classical induction proof that, in the wall convention, $v(x)=0$ if and only if $v(x+m)=1$, where $m=\min S$. We then explain how this can fail in the sink convention.
+Suppose $v(x)=0$ but $v(x+m)>1$.
+Then there exists a move $s$ such that v(x+ms)=1v(x+m-s)=1.
+Since $v(x)=0$, we have v(xs)>0v(x-s)>0 for all wall-feasible $s$.
+Hence subtracting $m$ from a heap of size x+msx+m-s does not yield a 0-position, which contradicts the induction hypothesis that every value “1” on a smaller heap must be $m$-preceded by a “0”.
+Conversely, suppose $v(x+m)=1$ but $v(x)>0$, which implies $v(x)>1$.
+Then there exists $s$ such that v(xs)=0v(x-s)=0.
+By induction, it follows that v(x+ms)=1v(x+m-s)=1, which contradicts the assumption that $v(x+m)=1$.
+In the sink convention, the base case fails; see for example Example 5, due to the presence of additional terminal zeroes.
+Observe that $v(5)=1$, but $v(3)\neq 0$.
+Indeed, the step “By induction, then v(x+ms)=1v(x+m-s)=1” does not apply, since in this case x+ms=0x+m-s=0, which belongs to the sink.
+Similarly, $v(8)=0$ but $v(10)\neq 1$.
+Here $v(8)$ is the first non-trivial zero-value, yet $v(10)\neq 1$ because $v(5)=1$.
+These violations of Ferguson’s principle propagate to larger heaps.
+Thus, in the sink convention, one can no longer deduce the $1$-values from the 0-values, or vice versa.
+Note that in the case of sink, the above conjecture, together with Ferguson’s pairing principle, instead implies a pairing of “1”s and “2”s: $v_{\mathrm{sink}}(x)=1$ if and only if $v_{\mathrm{sink}}(x-m)=2$.
 2 A proof of the main result
@@
 That is, if $W=(w_{i})_{i\geq 1}$ is the infinite word on the same alphabet such that, for all $i\geq 1$, $w_{i}$ is the nimber of $S$, then $W=PP\ldots$.
-Observation 6.
+Observation 7.
 In spite of the periodicity hinging on a quadratic expression in $m$, we are able to verify correctness in constant time.
@@
 It remains to justify the general case of item (ii), for a fixed $k$, and when $\delta=2nm+k$, for $n\geq 1$.
-For each increase in $n$, the period length increases by $4m^{2}/{\gcd(m,\delta)}$, and this is due to the fact that each block is increased by exactly $4m$ positions via the two concatenated factors $0^{m}1^{m}$ and $2^{m}3^{m}$; see Figure 4.
-The justification is similar to case (i), but here it invokes another quadratic term.
+For each increase in $n$, the period length increases by $4m^{2}/{\gcd(m,\delta)}$, and this is due to the fact that each block is increased by exactly $4m$ positions via the two concatenated factors $1^{m}2^{m}$ and $3^{m}0^{m}$.
+Precisely, for each increase in $n$, insert $1^{m}2^{m}$ just after each $A$-prefix and insert $3^{m}0^{m}$ exactly between the last zero-factor and the last three-factor (independently of whether a $B$- or $C$-block); see Figure 4.
+The mex-justification is similar to case (i), but here it invokes another quadratic term.
 Apart from this observation, the mex verification remains the same as for the case $n=1$.
@@
 The second part of the conjecture for three move games [3, 9] concerns the non-additive three-move rulesets.
-Although they all are conjectured to have period length “a sum of two moves”, and any such instance would individually be easily verified, the full classification appears to satisfy some fractal structure [3], and has not yet been fully understood.
+Although they all are conjectured to have period length “a sum of two moves”, and any such instance would individually be easily verified, the full classification appears to satisfy some fractal structure [3], which has not yet been fully understood.
 See also the recent survey on subtraction games [7].
+There is a closed formula expression for the $\mathcal{P}$-positions of additive wall subtraction in [2, 3]; and hence, by (1), it applies also to those of additive sink subtraction.
+Namely the set of wall $\mathcal{P}$-positions is generated by, for all $n\geq 0$,
+$n+\left\lfloor\frac{n}{m}\right\rfloor m+\left\lfloor\frac{2n}{2m+\delta-d}\right\rfloor(m+\delta)\ \text{ or }\ n+\left\lfloor\frac{n}{m}\right\rfloor m+\left\lfloor\frac{2n}{\delta+d}\right\rfloor(m+\delta),$
+depending on whether $d\leq m$ or not. To emphasize that the first (second) expression gives a linear (quadratic) period length, we rewrite this as
+$w_{n}=n+\left\lfloor\frac{n}{m}\right\rfloor m+\left\lfloor\frac{2n}{km}\right\rfloor(m+\delta)\ \text{ or }\ w_{n}=n+\left\lfloor\frac{n}{m}\right\rfloor m+\left\lfloor\frac{2n}{\ell m+2d}\right\rfloor(m+\delta),$
+with $k=(2m+\delta-d)/m\in\mathbb{N}$ and $\ell=(\delta-d)/m\in\mathbb{N}$, respectively.
+To illustrate, we get the following sequences for the parametrizations $m=2,\delta=d=1$ and $m=2,\delta=d=3$, respectively.
+$n$ 0 1 2 3 4 5 6 7 8 9 10 11
+$w_{n}(2,1)$ 0 1 7 8 14 15 21 22 28 29 35 36
+$s_{n}(2,1)$ 5 6 12 13 19 20 26 27 33 34 40 41
+$n$ 0 1 2 3 4 5 6 7 8 9 10 11
+$w_{n}(2,3)$ 0 1 4 10 13 14 22 23 26 32 35 36
+$s_{n}(2,3)$ 7 8 11 17 20 21 29 30 33 39 42 43
+Here, we have included, as a second data row, the corresponding sink-shifted $\mathcal{P}$-positions.
+Let us use the conjecture in Equation 3 to directly compute all the nimbers of both sink and wall for the first ruleset, $S=\{2,3,5\}$ using also Ferguson’s pairing principle.
+By the conjecture, from the $\mathcal{P}$-positions, we derive the nimber two positions for both conventions.
+For all $x\leq 12$, if $v_{w}(x)=0$, then $v_{s}(14-3-x)=2$. Hence we get the sink $2$-sequence: $3,4,10,11,\ldots$
+Similarly, for all $x\leq 17$, if $v_{s}(x)=0$, then $v_{w}(14+3-x)=2$. Hence we get the wall $2$-sequence: $4,5,11,12,\ldots$
+Next, the nimber one positions in the wall convention are obtained via the classic Ferguson’s pairing principle as $2,3,9,10,\ldots$.
+Now, the nimber three positions in the wall convention must be the remaining positions, namely: $6,13,20,\ldots$
+Similarly in the sink convention, the conjecture gives the postions of the nimber ones by the formula 143x14-3-x, where position $v_{w}(x)=1$.
+Thus, the sequence starts: $1,2,8,9,\ldots$ And the nimber three positions in the sink convention must be the remaining positions, namely: $5,12,19,\ldots$
+Note that all these values are obtained without any mex computation, from the classic $\mathcal{P}$-position formula in the wall convention.
+Similarly one can fill in the periodic nimber positions for the ruleset $\{2,5,7\}$, without using any mex-rule.
+Let us summarize these findings. Let $S=\{s_{1},s_{2},s_{3}\}$ be an additive ruleset, with $s_{1}<s_{2}<s3$.
+  1. 1.
+    By wall-sink outcome translation, $v_{w}(x)=0$ if and only if $v_{s}(x+s_{3}+1)=0$.
+  2. 2.
+    By Ferguson, wall zeros and ones are paired by the smallest move: $v_{w}(x)=0$ if and only $v_{w}(x+s_{1})=1$.
+  3. 3.
+    By our conjecture, given a wall value at position $x$, then $\sigma(v_{w}(x))=v_{s}(-s_{2}-x)$.
+  4. 4.
+    By our conjecture, given a sink value at position $x$, then $\sigma(v_{s}(x))=v_{w}(s_{2}-x)$.
+  5. 5.
+    Altogether by using the known formula of the $\mathcal{P}$-positions for the wall convention, we can therefore compute the positions of all nimbers in $\{0,1,2\}$ for both conventions.
+  6. 6.
+    Since the ruleset has exactly three moves, we may conclude that, for each convention, the remaining positions have value three, respectively.
+From our duality conjecture together with Ferguson’s pairing principle, let us derive explicit descriptions of all nimber classes in both conventions in terms of the wall $\mathcal{P}$-position sequence $(w_{n})$ and the period $p$ of the outcome patterns.
+For the wall convention we get the sets of positions with nimbers smaller than three as:
+$\mathcal{W}_{0}$ $=\{\,w_{n}:n\geq 0\,\},$
+$\mathcal{W}_{1}$ $=\{\,w_{n}+s_{1}:n\geq 0\,\},$
+$\mathcal{W}_{2}$ $=\{\,[\,-w_{n}-s_{1}\,]_{p}+kp:n\geq 0,\ k\geq 0\,\}.$
+And similarly for the sink convention:
+$\mathcal{S}_{0}$ $=\{\,w_{n}+s_{3}+1:n\geq 0\,\},$
+$\mathcal{S}_{1}$ $=\{\,[\,-s_{3}-w_{n}\,]_{p}+kp:n\geq 0,\ k\geq 0\,\},$
+$\mathcal{S}_{2}$ $=\{\,[\,-s_{2}-w_{n}\,]_{p}+kp:n\geq 0,\ k\geq 0\,\}.$
+As before, $x]_{p}$ denotes the smallest nonnegative representative of $x$ modulo $p$.
+In either case the nimber three positions are the remaining ones.
+Given our conjecture, in the case of “primitive quadratic” the nimber classes are also obtained from the wall bracket formulas [6] by linear transformations and reduction modulo $p$.
 Acknowledgements. We thank Richard Nowakowski for pointing out the reference [1]. The images in Figure 1 were generated with assistance from ChatGPT 5.2.